# Ditto: Motion-Space Diffusion for Controllable Realtime Talking Head Synthesis

### v2

@@ Abstract
Recent advances in diffusion models have revolutionized audio-driven talking head synthesis. Beyond precise lip synchronization, diffusion-based methods excel in generating subtle expressions and natural head movements that are well-aligned with the audio signal. However, these methods are confronted by slow inference speed, insufficient fine-grained control over facial motions, and occasional visual artifacts largely due to an implicit latent space derived from Variational Auto-Encoders (VAE), which prevent their adoption in realtime interaction applications. To address these issues, we introduce Ditto, a diffusion-based framework that enables controllable realtime talking head synthesis. Our key innovation lies in bridging motion generation and photorealistic neural rendering through an explicit identity-agnostic motion space, replacing conventional VAE representations. This design substantially reduces the complexity of diffusion learning while enabling precise control over the synthesized talking heads. We further propose an inference strategy that jointly optimizes three key components: audio feature extraction, motion generation, and video synthesis. This optimization enables streaming processing, realtime inference, and low first-frame delay, which are the functionalities crucial for interactive applications such as AI assistants. Extensive experimental results demonstrate that Ditto generates compelling talking head videos and substantially outperforms existing methods in both motion control and realtime performance.
@@ 1 Introduction
Audio-driven talking head synthesis [1, 2, 3, 4, 5, 6, 7] has emerged as a prominent research topic in recent years. In particular, one-shot methods [8, 9, 10, 11], which enable video generation from a single photograph, substantially reduce the data prerequisite for target identities, allowing for the creation of talking-head videos without relying on templates. In early research, approaches based on Generative Adversarial Networks (GANs) [12] are the predominant solution for this task. These studies demonstrate relatively accurate lip synchronization capabilities. However, they still lack diversity and realism in terms of speaking styles, emotions, and head movements. Recently, diffusion-based techniques [13, 14, 15], with EMO [3] as a notable example, have made substantial progress. These methods maintain excellent lip synchronization while further enhancing the liveliness and realism of generated videos. They capture subtle facial expressions, synchronize head movements with speech rhythm, and convey lifelike gaze interactions, which are crucial traits to achieve truly convincing synthesized results.
However, two critical issues hinder the broad application of these methods. The first issue is the lack of fine-grained control over facial motion. Existing approaches predominantly rely on audio and other weak conditions to influence the generated output, making it hard to enforce precise and delicate control over subtle aspects of a talking head, such as the driving regions, emotions and head movements. This limitation complicates the timely modification of small flaws in the generated results and the integration with undriven body regions, restricting applications in diverse scenarios. Moreover, small artifacts in teeth and hair are frequently observed in general text-to-image diffusions for portrait generation, further constraining the performance of these talking head approaches. The second issue is the relatively slow inference process. Currently, most diffusion-based approaches can hardly finish inference in realtime on a single GPU, which is a critical requirement for interactive scenarios. Therefore, although these methods offer enhanced vividness and expressiveness, the inference speed constraint largely prevents their deployment in realtime interactive applications such as AI assistants.
Looking into the aforementioned issues, we argue that their root cause is the use of a redundant, implicit, and inexplicable latent space for training the diffusion model. These approaches apply general text-to-image frameworks to the talking-head domain without decoupling the motion generation and photorealistic neural rendering stages [16]. Specifically, most current approaches adopt the Dual U-Net architecture introduced in EMO, training diffusion models within the latent space of a fixed pre-trained VAE. However, this generic latent space entangles motion and appearance representations, complicating both the diffusion process and subsequent decoding. While VASA-1 [17] achieved real-time inference by decoupling motion generation and facial rendering, it relies on an implicit facial representation, precluding explicit control over the generated motion. Nevertheless, this approach demonstrates the feasibility of integrating diffusion models with the representation space of facial reenactment.
We propose Ditto, a diffusion-based framework specifically designed for one-shot talking head synthesis. Ditto can produce lifelike videos from a single portrait image, enabling fine-grained facial motion control while achieving realtime inference. Our core idea is to leverage diffusion solely for identity-agnostic motion generation and employ a specialized one-shot face renderer for rapid, high-quality video synthesis. This design reduces the training complexity of the diffusion module and allows separate optimization of inference processes between modules.
Specifically, we define an identity-agnostic motion space that exclusively encompasses facial and head motions pertinent to talking-head animations. This space has a dimensionality ten times lower than conventional VAE spaces and establishes semantic correspondences with facial motions, thereby enabling fine-grained control over the generated results. We train a Diffusion Transformer (DiT) [18] in this space using audio, emotion labels, and eye states as conditions. Subsequently, we employ a dedicated one-shot decoder to render the generated motion features, achieving rapid and high-quality video synthesis.
Besides, we implement inference optimization suitable for streaming applications across three primary modules: audio feature extraction, DiT inference, and video synthesis. To our knowledge, our approach is the first to achieve low first-frame delay and realtime streaming inference for diffusion-based talking head generation. Our approach combines lifelike synthesis with rapid generation, providing foundational capabilities for innovative realtime dialogue applications such as AI assistants.
@@ 2 Related Work
In audio-driven talking head synthesis, one-shot methods have garnered research attention due to their ability to minimize data requirements to a single photograph. While early GAN-based approaches have succeeded in generating photorealistic textures and accurate lip movements, they fall short in capturing realistic expressions and head movements. Recent diffusion-based methods have made substantial progress in addressing these limitations, truly achieving the generation of realistic talking-head videos without relying on carefully curated templates. EMO [3] stands out as a groundbreaking approach that leverages audio and other “weak controls" to train a conditional diffusion model for end-to-end video generation, enhancing the vividness and realism of the generated results. Subsequent methods have widely adopted its innovative Dual U-Net architecture and two-stage training scheme. Notable follow-up works include EchoMimic [14], Hallo [15], and Loopy [13], which have improved the original framework in various aspects such as hierarchical audio-visual mapping, diverse driving signals, and long-term motion modeling, respectively. Nonetheless, these methods still rely on denoising and rendering images in a general VAE space. This redundant, implicit latent space not only increases the learning complexity of the Diffusion model but also results in slow inference speeds. VASA-1 [17] introduces a novel approach by training a Diffusion Transformer (DiT) [18] within a motion-appearance disentangled space, defined by a face reenactment model [19]. This latent space, tailored specifically for the facial domain, combined with a two-stage generation process, significantly reduces inference time while maintaining vibrant results, which has inspired our work. While, its use of implicit face representation lacks correspondence with facial semantics, hindering regional control and detail refinement in the generated results.
@@ 3 Method
Our approach enables the generation of realistic talking-head videos from a single portrait image and arbitrary speech audio clips, meanwhile supporting fine-grained control over facial attributes including emotions, headpose, and gaze. We begin by constructing a motion-appearance disentangled facial latent space, which serves as a bridge between motion generation and photorealistic rendering modules. Subsequently, we design a Diffusion Transformer (DiT) to learn the audio-to-motion mapping within this representation space. Finally, we present inference optimization for realtime deployment.
@@ 3.1 Motion Space
While Latent Diffusion Models (LDMs) [20] significantly reduce computational requirements by operating in latent space rather than pixel space, their pretrained VAE-defined latent representations present two key limitations in talking head synthesis. (1) The latent space remains redundant, with entangled motion and texture, introducing overhead in both diffusion training and pixel-level rendering. (2) These representations are implicit and lack explicit correspondence with facial attributes, limiting fine-grained control over the generation process. In the following sections, we demonstrate how our proposed motion space addresses these limitations.
Figure 2: Examples of dedicated fine-grained control of subtle facial motion. The notation “15-0" in the top-left image denotes the application of an x-axis offset to the 15th keypoint. The central image depicts the rendered result when applying zero offsets across all dimensions.
3.1.1 Disentangling Identity and Motion
Due to the shared underlying facial geometry, individuals across different identities exhibit highly consistent facial motion patterns. By constraining the diffusion model to learn only the mapping from audio to universal motions, while incorporating identity information during pixel generation, we can significantly reduce both the training and rendering complexity. The key to achieving this lies in the effective disentanglement of identity and motion. Inspired by recent facial reenactment works [21, 22, 23, 24, 25, 26], we construct an explicit motion space based on expression deformation and head rotation.
@@ 3.1.1 Disentangling Identity and Motion
Then we devise a one-shot face renderer comprising an Appearance Feature Extractor $\mathcal{F}$, a Face Renderer $\mathcal{G}$ [27] to synthesize talking head videos. Given a target portrait, $\mathcal{M}$ first extracts its canonical keypoints $\mathbf{c_{ref}}$ and motion representation $\mathbf{m_{ref}}=\{\mathbf{\boldsymbol{\delta}_{ref}},\mathbf{R_{ref}},\mathbf{t_{ref}}\}$. Then the reference and generated implicit 3D keypoints ($\mathbf{x_{ref}}$ and $\mathbf{\hat{x}}$) are computed as:
@@
3.1.2 Motion Control
Unlike headpose which can be directly controlled through rotation and translation, 3D implicit keypoint-based motion representations [21, 22, 23] lack clear correspondence with facial attributes. To address this limitation, [27] employs landmark-guided optimization to align compact implicit keypoints with explicit landmarks. We extend this work by establishing the direct mapping between deformations and facial semantics, enabling motion control in a blendshape-based manner. Specifically, we represent expression deformation as a 63-D vector, corresponding to the x, y, and z coordinates of 21 implicit 3D keypoints. For each dimension, we apply small positive and negative offsets and render facial images for a fixed identity. As illustrated in Figure 2, each dimension of deformations affects relatively distinct facial regions, generating corresponding local variations. For instance, the 34th dimension governs the right eye’s opening and closing motion. The 58th dimension controls mouth open, analogous to the “jaw open" blendshape in ARKit. Therefore, we enable two types of generation control: (1) Regional control, where motion generation can be constrained to specific local facial areas. For example, limiting movements to facial features close to the neck for better integration with the torso, and (2) Magnitude control, where we can impose constraints on deformation values to prevent artifacts such as unnatural facial expressions.
Figure 3: Visualization of how ICS (Initial Conditional Signal) and ECS (Enhanced Conditional Signal) are fed into our DiT-based Motion Generator.
@@ 3.2 Motion Generation with Diffusion Transformer
Consistent with prior works [13, 15, 14], we employ diffusion models for audio-conditioned video generation. Differently, we eschew the Dual-Unet structure [3] and two-stage training in both spatial and temporal stages. Instead, we utilize a transformer network [28], leveraging the compact motion representations, to accomplish the motion generation in time series.
3.2.1 Conditional Information
As shown in Figure 3, we employ a DiT for audio-to-motion generation. Along with the audio feature $\mathbf{a}$, we incorporate multiple auxiliary conditional signals as follows. We use the canonical keypoints as the identity feature $\mathbf{c_{ref}}$ to guide the generated facial motions, ensuring compatibility with the target’s facial structure. This identity-adaptive setting mitigates the long-standing issue of incomplete disentanglement between motion and identity [29, 30]. Using a visual emotion classifier [31], we extract emotions from video frames and employ a unique emotion label $\mathbf{s}$ at the clip level, establishing more direct and controllable correspondence compared to traditional audio-to-expression mapping. Eye state $\mathbf{e}$ is composed of the aspect ratio and relative pupil position [32], governing blinking and gaze which have weak correlations with audio. Dedicated conditions are beneficial for generating accurate results. A reference initial motion $\mathbf{m_{ref}}$ serves as a guide for the motion generation within each clip, which can be utilized to enhance inter-clip motion continuity, as well as to periodically steer the motion towards a specified target state, thereby mitigating error accumulation in long sequence generation [33].
@@ 3.2.2 Training Strategy
We devise specialized strategies for training the DiT to accommodate the characteristics of the talking head scenario and the proposed motion representation.
@@
Adaptive Loss Weights: The correlation between audio and lip movements is highly consistent across different identities. However, the relationship between audio and facial expression is quite complicated, furthermore, the correlation between audio and headpose is highly individualistic. It is hard to train an audio-driven model by uniformly weighing different motion representations. Therefore, we group the motion representations and dynamically adjust the weights of each group throughout the training process. Experiments demonstrate that this dynamic weighting strategy improves both convergence speed and the quality of generated motion.
@@
3.2.3 Loss Function
Following DDPM [34], a denoising network $\mathcal{T}$ is trained by minimizing the mean square error loss as follows:
@@ 3.2.3 Loss Function
Table 1: Comparisons with existing methods. The s* represents the number of denoising steps. The best and second best results are in bold and underline specifically.
@@ 3.3.1 Audio Feature
Figure 4: The qualitative comparison of our approach on two characters with different styles, poses, and hairstyles. For each character, we generat videos using the same audio input with each method and select frames at the same location for comparison. Blue arrows indicate the locations of artifacts, including inaccurate lip movements, disordered teeth, blur, etc.
@@ 4.1 Implementation Details
For training data, we collect and clean about 50 hours of broadcast scene videos from 330 identities, with an average video length of 150 seconds. For testing, we use the videos collected by previous works [39, 40, 41, 42] for quantitative evaluations. To evaluate the generalizability of our method, we select diverse image and audio samples from [15, 14] for qualitative evaluations.

### v3

@@ Abstract
Recent advances in diffusion models have endowed talking head synthesis with subtle expressions and vivid head movements, but have also led to slow inference speed and insufficient control over generated results. To address these issues, we propose Ditto, a diffusion-based talking head framework that enables fine-grained controls and real-time inference. Specifically, we utilize an off-the-shelf motion extractor and devise a diffusion transformer to generate representations in a specific motion space. We optimize the model architecture and training strategy to address the issues in generating motion representations, including insufficient disentanglement between motion and identity, and large internal discrepancies within the representation. Besides, we employ diverse conditional signals while establishing a mapping between motion representation and facial semantics, enabling control over the generation process and correction of the results. Moreover, we jointly optimize the holistic framework to enable streaming processing, real-time inference, and low first-frame delay, offering functionalities crucial for interactive applications such as AI assistants. Extensive experimental results demonstrate that Ditto generates compelling talking head videos and exhibits superiority in both controllability and real-time performance.
@@ 1 Introduction
Audio-driven talking head synthesis [1, 2, 3, 4, 5, 6, 7] has emerged as a prominent research topic in recent years. In particular, one-shot methods [8, 9, 10, 11] substantially reduce the data prerequisite for target identities, no longer relying on templates. Early approaches based on Generative Adversarial Networks (GANs) [12] are the predominant solutions, showing accurate lip synchronization but still lacking diversity and realism. Recently, diffusion-based techniques [13, 14, 15], with EMO [3] as a notable example, have made substantial progress. These methods maintain lip synchronization and capture subtle facial expressions, synchronizing head movements with speech rhythm, which are crucial traits to achieve compelling synthesized results.
However, two critical issues hinder the broad application of these diffusion-based methods. (1) lack of control over facial motion. Existing approaches struggle to enforce fine-grained control over talking heads, including facial movements, basic emotions, and head rotations. This leaves users with no direct means to adjust the results in addition to regeneration. As the generation quality is relatively random, achieving desired results becomes challenging and potentially time-consuming. (2) Slow inference speed. Currently, most methods can hardly achieve real-time inference on a single GPU, which is a critical requirement for interactive scenarios, such as AI assistants and live video streaming. VASA-1 [16] innovatively achieves real-time inference through a two-stage generation approach, demonstrating the feasibility of training a Diffusion Transformer (DiT) [17] based on a motion space. However, due to its source code is not publicly available, related follow-up work is limited. On the other hand, VASA-1 uses an implicit motion representation, which does not support control or adjustment of the generated results.
In this paper, we propose Ditto, a diffusion-based talking head framework that enables fine-grained controls and real-time inference. Specifically, we utilize an off-the-shelf motion extractor [18] and devise a conditional diffusion transformer to generate representations in a specific motion space. We optimize the model architecture and training strategy to address inherent issues in motion representations, including insufficient disentanglement between motion and identity information, and large internal discrepancies within the motion representations.
Besides, we employ diverse conditional signals to achieve fine-grained control over the generation process. We also establish a mapping between motion representation and facial semantics, enabling fine-tuning of facial movements and correction of visual defects.
Moreover, we jointly optimize the holistic framework to enable streaming processing, real-time inference, and low first-frame delay, offering functionalities crucial for interactive applications such as AI assistants. Extensive experimental results demonstrate that Ditto generates compelling talking head videos and exhibits superiority in both controllability and real-time performance. For the advancement of the community, we make our source code available for open-source use.
@@ 2 Related Work
In audio-driven talking head synthesis, one-shot methods have garnered research attention due to their ability to minimize data requirements to a single photograph. While early GAN-based approaches have succeeded in generating photorealistic textures and accurate lip movements, they fall short in capturing realistic expressions and head movements. Recent diffusion-based methods have made substantial progress in addressing these limitations, truly achieving the generation of realistic talking-head videos without relying on carefully curated templates. EMO [3] stands out as a groundbreaking approach that leverages audio and other “weak controls” to train a conditional diffusion model for end-to-end video generation, enhancing the vividness and realism of the generated results. Subsequent methods have widely adopted its innovative Dual U-Net architecture and two-stage training scheme. Notable follow-up works include EchoMimic [14], Hallo [15], and Loopy [13], which have improved the original framework in various aspects such as hierarchical audio-visual mapping, diverse driving signals, and long-term motion modeling, respectively. Nevertheless, these methods still rely on denoising and rendering images in a general VAE space. This redundant, implicit latent space not only increases the learning complexity of the Diffusion model but also results in slow inference speeds. VASA-1 [16] introduces a novel approach by training a Diffusion Transformer (DiT) [17] within a motion-appearance disentangled space, defined by a face reenactment model [19]. This latent space, tailored specifically for the facial domain, combined with a two-stage generation process, significantly reduces inference time while maintaining vibrant results.
@@ 3 Method
In this section, we introduce the detailed implementation of the Ditto framework. We begin with a brief introduction to the motion space. Then we discuss the model architecture and training strategy to address inherent issues in motion representations. Finally, we introduce a novel motion control approach and demonstrate how to correct visual defects using gaze as an example.
@@ 3.1 Motion Space
While Latent Diffusion Models (LDMs) [20] significantly reduce computational requirements by operating in latent space rather than pixel space, their pretrained VAE-defined latent representations present two key limitations in talking head synthesis. (1) The latent space remains redundant, with entangled motion and texture, introducing overhead in both diffusion training and pixel-level rendering. (2) These representations are implicit and lack explicit correspondence with facial attributes.
Due to the shared underlying facial geometry, individuals across different identities exhibit highly consistent facial motion patterns. Recent facial reenactment works [21, 22, 23, 24, 25, 26] also fully utilize this characteristic. We construct the motion space based on LivePortrait [18], one of the state-of-the-art methods in this field. We aim to let the diffusion model generate universal motion, and then incorporate identity information during pixel-level rendering.
@@
Then a one-shot face renderer, comprising an Appearance Feature Extractor $\mathcal{F}$ and a Face Renderer $\mathcal{G}$ [18], is used to synthesize talking head videos. Given a target portrait, $\mathcal{M}$ first extracts its canonical keypoints $\mathbf{c_{ref}}$ and motion representation $\mathbf{m_{ref}}=\{\mathbf{\boldsymbol{\delta}_{ref}},\mathbf{R_{ref}},\mathbf{t_{ref}}\}$. Then the reference and generated implicit 3D keypoints ($\mathbf{x_{ref}}$ and $\mathbf{\hat{x}}$) are computed as:
@@ 3.2 Motion Generation with Diffusion Transformer
As shown in Figure 2, we employ a Conditional Diffusion Transformer (DiT) for audio-to-motion generation. Along with the audio feature $\mathbf{a}$, we incorporate multiple auxiliary conditional signals as follows.
3.2.1 Conditional Signals
Considering the fact that even LivePortrait [18] is hard to fully disentangle motion and identity information, and the impure motion representations increase the difficulty of the training process. We use the canonical keypoints as the identity feature $\mathbf{c_{ref}}$ to guide the generation of facial motions, ensuring the generated motions match the target facial characteristics and can appropriately animate the target identity. This identity-adaptive setting mitigates the long-standing issue of incomplete disentanglement between motion and identity [27, 28]. Ablation experiments also confirm the above conclusions.
We also extract emotions [29] from video frames and employ a unique emotion label $\mathbf{s}$ at the clip level, establishing a direct and controllable correspondence instead of simply learning the audio-to-expression mapping. Eye state $\mathbf{e}$ is composed of the aspect ratio and relative pupil position [30], governing blinking and gaze which have weak correlations with audio. A reference initial motion $\mathbf{m_{ref}}$ serves as a guide for the motion generation within each clip, which can be utilized to enhance inter-clip motion continuity, as well as to periodically steer the motion towards a specified target state, thereby mitigating error accumulation in long sequence generation [31].
@@ 3.2.1 Conditional Signals
Figure 2: Visualization of the architecture of the proposed conditional Diffusion Transformer, which generates compact motion representations based on various conditional signals.
@@ 3.2.2 Training Strategy
We devise specialized strategies for training the DiT to accommodate the specific motion representations.
@@
Adaptive Loss Weights: Different components of the motion representation exhibit variations in their movement patterns. For example, motions such as lip movements, facial expressions, eye blinking, and head pose differ significantly in their relationships with driving signals like audio, and in the magnitude of their movements. It is difficult to train the DiT to weigh all motion representations uniformly. Therefore, we group the motion representations based on their control regions, and dynamically adjust the weights of each group throughout the training process. We calculate the difference in average loss between the current epoch and the previous epoch, then apply softmax normalization across groups to derive adjustment coefficients for the loss weights. Experiments demonstrate that this dynamic weighting strategy improves both the loss convergence and the quality of generated motion.
Figure 3: Examples of dedicated fine-grained control of subtle facial motion. The notation “15-0” in the top-left image denotes the application of an x-axis offset to the 15th keypoint. The central image depicts the rendered result when applying zero offsets across all dimensions.
@@
Loss Function: Following DDPM [32], a denoising network $\mathcal{T}$ is trained by minimizing the mean square error loss as follows:
@@
3.2.3 Motion Control
Unlike headpose which can be directly controlled via rotation and translation, our implicit keypoint-based motion representations [21, 22, 23] lack clear correspondence with facial attributes. To address this limitation, [18] employs landmark-guided optimization to align implicit keypoints with explicit landmarks. We extend this work by establishing the direct mapping between deformations and facial semantics, enabling motion control in a blendshape-like manner. Specifically, we represent expression deformation as a 63-D vector, corresponding to the x, y, and z coordinates of 21 implicit 3D keypoints. For each dimension, we apply small positive and negative offsets and render facial images for a fixed identity. As illustrated in Figure 3, each dimension of deformations affects relatively distinct facial regions. For instance, the 34th dimension governs the right eye’s opening and closing motion. The 58th dimension controls mouth open, analogous to the “jaw open” blendshape in ARKit. Therefore, we enable two types of generation control: (1) Regional control, where motion generation can be constrained to specific local facial areas. For example, limiting movements to facial features close to the neck for better integration with the torso, and (2) Magnitude control, where we can impose constraints on deformation values to prevent artifacts such as unnatural facial expressions.
Next, we introduce how to correct visual defects in the generated motions through a case of gaze adjustment. In the early test, we discover that the generated gaze is bound to the head pose. This results in the avatar’s gaze constantly wandering with head movements in the generated video, unable to focus on the camera. The reason is that we control the gaze by adding the per frame eye state $\mathbf{e}$ during training. But we use the fixed eye state during inference, which is extracted from the single reference image.
To address this issue, we record a template video, requiring the actor to smoothly rotate head in various angles while maintaining eye focus on the camera and a constant facial expression. Then we extract the corresponding head poses $\mathbf{R}$ and expression deformations $\boldsymbol{\delta}$. By subtracting the initial frame’s $\mathbf{R_{0}}$ and $\boldsymbol{\delta_{0}}$ from the $\mathbf{R}$ and $\boldsymbol{\delta}$ of each frame, we obtain synchronized variation data for pose and motion. Since the actor’s expression remains constant throughout the video, the motion variation only reflects the change in gaze. We use a regression $\boldsymbol{\delta_{e}}=\mathcal{K}(\boldsymbol{R_{e}})$ to obtain the mapping between the gaze variation $\boldsymbol{\delta_{e}}$ and the head pose variation $\boldsymbol{R_{e}}$, which is used to correct the generated gaze based on the current headpose, $\boldsymbol{\delta}_{correct}=\boldsymbol{\delta}+\boldsymbol{\delta_{e}}$.
Table 1: Comparisons with existing methods on the Talk9 dataset. The s* represents the number of denoising steps. The best and second best results are in bold and underline specifically.
@@ 3.2.3 Motion Control
[…]
Method FID $\downarrow$ FVD $\downarrow$ SyncC $\uparrow$ SyncD $\downarrow$
EchoMimic 47.331 532.733 5.930 9.143
Hallo 16.748 366.066 7.268 7.714
Hallo2 16.616 239.517 7.379 7.697
Ours 16.430 134.640 8.939 6.913
@@ 4.1 Implementation Details
For training data, we collect and clean about 50 hours of broadcast scene videos from 330 identities, with an average video length of 150 seconds. For evaluation, due to the limited computing resources, we conduct our primary quantitative comparison on Talk9 dataset while providing additional quantitative comparison on HDTF100 dataset. The Talk9 dataset consists of 9 videos from [37, 38, 39, 40] and has been widely used in GAN/NeRF-based talking head evaluations. HDTF100 contains 100 randomly sampled videos from the HDTF [41] dataset, which has been widely used for quantitative evaluation in previous works [14, 15, 42]. To evaluate the generalizability of our method, we select diverse image and audio samples from [15, 14] for qualitative evaluations